Over a period of time Julian finds that on long-distance flights he flies economy class on 82% of flights. On the rest of the flights he flies first class. When he flies economy class, the probability that he gets a good night's sleep is x. When he flies first class, the probability that he gets a good night's sleep is 0.9.
- (i)Draw a fully labelled tree diagram to illustrate this situation.[2]
- (ii)The probability that Julian gets a good night's sleep on a randomly chosen flight is 0.285. Find the value of x.[2]
- (iii)Given that on a particular flight Julian does not get a good night's sleep, find the probability that he is flying economy class.[3]
This question uses a diagram from the original paper, which isn't shown here. Open the full paper from the past papers archive to see it.
Mark scheme
Marking points from the official mark scheme (Oct/Nov 2017). The mark or mark type is shown in brackets.
- [1] (i) Must see at least 4 probs correct including one with an x in, correct shape
- [1] (i) Shape, clear labels/annotation and all probs correct
- [1] (ii) 0.82x + 0.18 × 0.9 = 0.285 (Eqn with x in, two 2-factors on one side)
- [1] (ii) x = 0.15
- [1] (iii) Attempt at P(E∩not GNS) seen as num or denom of fraction
- [1] (iii) Attempt at P(not GNS) seen anywhere
- [1] (iii) Correct answer = 0.975
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →